OpenStudy (anonymous):

what is the average(arithmetic mean) of the positive integers from 1 to 50 inclusive?

OpenStudy (anonymous):

add up all the number between 1 and 50 (including 1 and 50) and then divide that amount by how many digits are between 1 and 50(inculsive)

OpenStudy (anonymous):

Isn't there a shortcut?

OpenStudy (anonymous):

this is a PSAT question by the way

OpenStudy (anonymous):

here is what you do: google: shortcut for arithmitec mean

OpenStudy (anonymous):

after you do that, go play pokemon

OpenStudy (anonymous):

Think about it logically: 1 + 50 = 51 2 + 49 = 51 3 + 48 = 51 ... 25 + 26 = 51 These 25 pairs of numbers all equal 51 - If you add them all together, you'd be multiplying 51 by 25, then dividing by 50 to get the average. So, you're multiplying 51 by (25/50) or taking: \[\frac{1 }{ 2 }*50\] The answer is 25.5.